The bulldozer exerts a force F = 2i (kip) at A. wnat are the tensions in cables AB, AC, and AD? ( Poi 6式 8 ft 4 ft 8 fte
Answer #1

Pocii on vectors s AC = (0,276 )-(8,010) AB = (0,3,8)-(8,0,0) AG 8+3 k. NOU AC AC = AC AC V 6 +4+36TAA ARl AB 6u4+64 AD TAD-tan ADI 」 6$- 43 6u tl6 Free pody nicgam AC AB TAD -0.184 TAC-O, 68 3 TAB-o.894TA D=-2000256RE o.uu7%b-0 -2 -o.$88 TAC 0.683T $88 TAC .683 TAB= o. 860 TAC AD. ADO.IC T o.4un A AD.30 TAC2000 AC- 2-20 TA B = 0.960 (907.93)